2X-4=X over 2-3
what is X ?

Answer by stanbon(48535)

2X-4=X over 2-3
2x-4 = (x/2)-3
2x=(x/2)+1
2x-(1/2)x=1
(3/2)x=1
x=2/3
Cheers,
Stan H.